OpenStudy (anonymous):

A (1+ T/100) (1+ T/100)=

OpenStudy (anonymous):

\[A \left(1+\frac{T}{100}\right)^2=A \left(1+\frac{T}{50}+\frac{T^2}{10000}\right)=\frac{A \left(10000+200 T+T^2\right)}{10000} \]